# SmartQuant: CXL-based AI Model Store in Support of Runtime Configurable Weight Quantization

### v1

@@ I Introduction
With the wide acceptance of the scaling law [1], AI model size will continue to increase at least in the foreseeable future, despite today’s AI models already contain over a trillion parameters [2, 3]. To unleash the full potential of large AI models in the real world, it is essential to maximize the inference cost/energy efficiency, for which quantization is one of the most effective means and has been widely studied [4, 5, 6]. To materialize the benefits of quantization, modern GPU and AI accelerators provide strong built-in hardware support for variable-precision arithmetics. For modern generative AI models such as transformer [7], recent studies show that the importance of different weights exhibits substantial context-dependent variations during inference [8]. This observation opens new opportunities to improve the inference efficiency via adaptive weight quantization configuration. Although runtime weight quantization configuration can leverage the hardware support for variable-precision arithmetics in GPU and AI accelerators, little prior research has studied how one could exploit configurable weight quantization to adaptively reduce the AI model load latency and energy consumption.
@@
To fill this missing link, this letter presents a CXL-based AI model store that can gracefully support runtime configurable weight quantization. Built upon the rapidly maturing CXL ecosystem with native support of memory pooling/sharing [9], CXL memory devices are well positioned to play a key role in future computing infrastructure. Since today’s CPUs already have a large number of PCIe lanes (e.g., AMD Genoa has 128 PCIe Gen5 lanes), it is reasonable to expect future CPU/GPU and AI accelerators could readily support 150+ PCIe lanes. Upcoming PCIe Gen7 specifies a per-lane bandwidth of 16GB/s, leading to total 4TB/s bandwidth (2TB/s in each direction) with 128 PCIe lanes. In comparison, the upcoming 4th generation HMB (high bandwidth memory) is expected to enable $\sim$1.5TB/s throughput per DRAM stack, and one GPU package may integrate 4$\sim$8 HBM DRAM stacks. Hence, CXL memory devices could be a viable cost-effective supplement (or even alternative) to HBM for building future AI systems. This letter advocates for a CXL-based AI model store that can serve generative AI inference with dynamically configurable weight quantization (e.g., FP16, FP8, FP6). Its practical implementation is subject to two major issues: (1) How to materialize the potential of configurable weight quantization on improving the AI model DRAM access energy efficiency and throughput? As the straightforward design option, CXL memory controller reads the full-precision weights from DRAM and truncates/converts them into lower-precision ones, which however will not reduce the model load latency and energy consumption at all. (2) How to simplify the system integration? Inference computing devices should be able to conveniently configure and load AI model weights from CXL memory devices. To address these issues, this letter presents a design solution called SmartQuant that consists of two simple yet effective methods, including bit-plane in-memory placement and memory logical space bloating. Using the open-source transformer model [3] as a test vehicle and DRAM simulator DRAMSim3 [10], we carried out experiments to study the effectiveness of the proposed design solution. In comparison to the straightforward design, the results show that, SmartQuant can reduce the model load latency by up to 42.1% and memory access energy consumption by 40.3%.
@@ III-A Bit-Plane In-Memory Placement
To address this issue, the only viable option is to bit-wise disaggregate the in-memory storage of each full-precision weight, similar to a bite-wise column-store. Let $N_{1}$ denote the number of bits in full-precision weight quantization. We store all the model weights in $N_{1}$ bit-planes, each bit-plane corresponding to one bit position in the $N_{1}$-bit full-precision quantization. All the $N_{1}$ bit-planes are stored in DRAM independetly from each other, as illustrated in Fig. 2. Through such bit-plane in-memory placement, CXL memory controller can selectively fetch a just-enough portion of each full-precision weight from DRAM to realize quantization format conversion. Suppose the full-precision quantization contains 1-bit sign, $n_{e}$-bit exponent, and $n_{m}$-bit mantissa (i.e., $N_{1}=1+n_{e}+n_{m}$). Assume inference computing devices request a chunk of weights under a reduced-precision quantization with 1-bit sign, $r_{e}$-bit exponent, and $r_{m}$-bit mantissa. We could perform quantization format conversion by fetching data from $1+(r_{e}+d_{e})+(r_{m}+d_{m})$ in-memory bit-planes, where $d_{e},d_{m}\geq 0$. When setting $d_{e}=d_{m}=0$, we obtain the reduced-precision model weights by directly truncating the full-precision model weights. When setting $d_{e}$ and/or $d_{m}$ as 1 or 2, we can improve the quantization format conversion accuracy by rounding-off the full-precision model weights. Utilizing such bit-plane in-memory weight placement, we could proportionally reduce the amount of DRAM page activation and data transfer, leading to the proportional DRAM access efficiency.

### v2

@@ I Introduction
With the wide acceptance of the scaling law [1, 2], AI model size will continue to increase at least in the foreseeable future, despite today’s AI models already contain over a trillion parameters [3, 4]. To unleash the full potential of large AI models in the real world, it is essential to maximize the inference cost/energy efficiency, for which quantization is one of the most effective means and has been widely studied. To materialize the benefits of quantization, modern GPU and AI accelerators provide strong built-in hardware support for variable-precision arithmetics. For modern generative AI models such as transformer [2], recent studies show that the importance of different weights exhibits substantial context-dependent variations during inference [5]. This observation opens new opportunities to improve the inference efficiency via adaptive weight quantization configuration. Although runtime weight quantization configuration can leverage the hardware support for variable-precision arithmetics in GPU and AI accelerators, little prior research has studied how one could exploit configurable weight quantization to adaptively reduce the AI model load latency and energy consumption.
@@
To fill this missing link, this letter presents a CXL-based AI model store that can gracefully support runtime configurable weight quantization. Built upon the rapidly maturing Compute Express Link (CXL) ecosystem with native support of memory pooling/sharing [6], CXL memory devices are well positioned to play a key role in future computing infrastructure. Since today’s CPUs already have a large number of PCIe lanes (e.g., AMD Genoa has 128 PCIe Gen5 lanes), it is reasonable to expect future CPU/GPU and AI accelerators could readily support 150+ PCIe lanes. Upcoming PCIe Gen7 specifies a per-lane bandwidth of 16GB/s, leading to total 4TB/s bandwidth (2TB/s in each direction) with 128 PCIe lanes. In comparison, the upcoming 4th generation HBM (high bandwidth memory) is expected to enable $\sim$1.5TB/s throughput per DRAM stack, and one GPU package may integrate 4$\sim$8 HBM DRAM stacks. Hence, CXL memory devices could be a viable cost-effective supplement (or even alternative) to HBM for building future AI systems. This letter advocates for a CXL-based AI model store that can serve generative AI inference with dynamically configurable weight quantization (e.g., FP16, FP8, FP6). Its practical implementation is subject to two major issues: (1) How to materialize the potential of configurable weight quantization on improving the AI model DRAM access energy efficiency and throughput? As the straightforward design option, CXL memory controller reads the full-precision weights from DRAM and truncates/converts them into lower-precision ones, which however will not reduce the model load latency and energy consumption at all. (2) How to simplify the system integration? Inference computing devices should be able to conveniently configure and load AI model weights from CXL memory devices. To address these issues, this letter presents a design solution called SmartQuant that consists of two simple yet effective methods, including bit-plane in-memory placement and memory logical space bloating. Using the open-source transformer model [4] as a test vehicle and DRAM simulator DRAMSim3 [7], we carried out experiments to study the effectiveness of the proposed design solution. In comparison to the straightforward design, the results show that, SmartQuant can reduce the model load latency by up to 42.1% and memory access energy consumption by 40.3%.
@@ II Background and Motivation
Compute Express Link (CXL) is a high-speed interconnect standard that enhances communication between computing engines and memory. Built on PCIe, CXL adds protocols for low-latency, high-bandwidth data transfer, crucial for AI and data-intensive tasks. It includes CXL.io for I/O, CXL.cache for caching in accelerators, and CXL.mem for flexible memory access. CXL’s customizability and memory pooling support make it ideal for meeting the dynamic demands of modern computing [6].
@@ III-A Bit-Plane In-Memory Placement
To address this issue, the only viable option is to bit-wise disaggregate the in-memory storage of each full-precision weight, similar to a bit-wise column-store. Let $N_{1}$ denote the number of bits in full-precision weight quantization. We store all the model weights in $N_{1}$ bit-planes, each bit-plane corresponding to one bit position in the $N_{1}$-bit full-precision quantization. All the $N_{1}$ bit-planes are stored in DRAM independently from each other, as illustrated in Fig. 2. Through such bit-plane in-memory placement, CXL memory controller can selectively fetch a just-enough portion of each chunked full-precision weights from DRAM to realize quantization format conversion. Suppose the full-precision quantization contains 1-bit sign, $n_{e}$-bit exponent, and $n_{m}$-bit mantissa (i.e., $N_{1}=1+n_{e}+n_{m}$). Assume inference computing devices request a large chunk of weights under a reduced-precision quantization with 1-bit sign, $r_{e}$-bit exponent, and $r_{m}$-bit mantissa. We could perform quantization format conversion by fetching data from $1+(r_{e}+d_{e})+(r_{m}+d_{m})$ in-memory bit-planes, where $d_{e},d_{m}\geq 0$. When setting $d_{e}=d_{m}=0$, we obtain the reduced-precision model weights by directly truncating the full-precision model weights. When setting $d_{e}$ and/or $d_{m}$ as 1 or 2, we can improve the quantization format conversion accuracy by rounding-off the full-precision model weights. Utilizing such bit-plane in-memory weight placement, we could proportionally reduce the amount of DRAM page activation and data transfer, leading to the proportional DRAM access efficiency.